if the discount rate is 6%, what is the present value of a project that is estimated to yeild quarterly cash flows of $2900 for the next 21 years?
Answer #1

Quarterly discount rate = 6%/4 = 1.5% = 0.015

No. of cashflows = 21*4 = 84

So, Present value = 2900/1.015 + 2900/1.015^2 + ... + 2900/1.015^84

=2900/0.015*(1-1/1.015^84) from the PV of annuity formula

=$137978.04

  

Assumptions - The discount rate is a stated rate and Cashflows occur at the end of each quarter